OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 24 inches and a height of 16 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (anonymous):

V(A) = 12^2*8*1/3 V(A)=383.96 V(B) = 24^2*16*1/3 V(B) = 3,071.97 3,071.97/383.96= 8

OpenStudy (anonymous):

V(A) is 8 times smaller than v(b) Though I suggest you round answers to 384 and 3072

OpenStudy (anonymous):

Thank you! I finally figured it out after I looked at it again!!

OpenStudy (anonymous):

@ciarriacuster Sweet!! Glad i could help :)